Lucido Products markets two computer games: Claimjumper and Makeover. A contribution format income statement for a recent month for the two games appears below:

Claimjumper Makeover Total
Sales $ 116,000 $ 58,000 $ 174,000
Variable expenses 35,800 7,700 43,500
Contribution margin $ 80,200 $ 50,300 130,500
Fixed expenses 87,750
Net operating income $ 42,750

Required:

1. What is the overall contribution margin (CM) ratio for the company?

2. What is the company's overall break-even point in dollar sales?

3. Prepare a contribution format income statement at the company's break-even point that shows the appropriate levels of sales for the two products.
Answer #1

Requirement 1

Overall Contribution margin ratio 75%

Requirement 2

Overall Break-even point $ 117,000.00

Requirement 3

Lucido Products
Contribution income statement
Claimjumper Makeover Total
Sales $          78,000 $              39,000 $        117,000
Variable expenses $          24,072 $                 5,178 $           29,250
Contribution margin $          53,928 $              33,822 $           87,750
Fixed cost $           87,750
Net operating income (loss) $                    -  

Working

Total
Sales $ 174,000.00
Variable expenses $    43,500.00
Contribution margin $ 130,500.00
Contribution margin ratio (130500/174000) 75%
Fixed cost $    87,750.00
Overall Break-even point (87750/75%) $ 117,000.00

.

Lucido Products
Contribution income statement
Claimjumper Makeover
Sales =116000/174000*117000 =58000/174000*117000
Variable expenses =35800/116000*78000 =7700/58000*39000